# Learning Encodings by Maximizing State Distinguishability: Variational Quantum Error Correction

### v1

@@ 4.3 Generating Recovery Operations
While the main focus of our paper is the encoding operation, it is also important to construct fitting recovery operations $U_{\mathrm{rec}}$ for practical error correction. Therefore, we adopt a learning-based method inspired by the QVECTOR procedure [46] to obtain optimal parameters $\Phi$ for a variational ansatz $U_{\mathrm{rec}}(\Phi)$. The details and adaptations we made to the original procedure are deferred to Appendix C. In the following, we provide just a superficial overview:
@@
$\mathcal{F}_{\mathcal{S}}(\mathcal{N};\Theta,\Phi)=1-\frac{1}{\left|\mathcal{S}\right|}\sum_{\rho\in\mathcal{S}}F(\rho,\hat{\rho}),$ (21)
@@ 4.4 Properties of Learned Codes
However, we find that certain instances of VarQEC codes can achieve exact error correction. For example, in Section B.3 we show that VarQEC codes with parameters [[$n\geq 5,1$]] trained on depolarizing noise exhibit a code distance of $d=3$, matching the [[$5,1,3$]] perfect code [53]. This analysis is conducted usind a novel empirical method we introduce in Section D.1: Based on the worst-case distinguishability loss, it determines the potential code distance of any QEC or VarQEC code. However, we again want to emphasize that the objective of VarQEC is not to generate codes with a certain provable code distance, but to develop low-overhead codes tailored for the specific noise channel.
As further motivation for training with the distinguishability loss, we derived a theoretical relationship to the worst-case fidelity after recovery. We state this connection informally here, with the details provided in Section B.2. It is important to note that certain mild assumptions on the noise channel are required, which are, however, satisfied for typical channels like depolarizing noise or bit-flip noise with a strength of $p\leq\frac{2}{3}$.
@@
The worst-case distinguishability loss $\overline{\mathcal{D}}$ under a well-behaved noise channel $\mathcal{N}$ guarantees the existence of a recovery operation that achieves a worst-case fidelity loss bounded by:
¯$\overline{\mathcal{F}}(\mathcal{N})\leq 1-\left(1-\overline{\mathcal{D}}(\mathcal{N})\right)^{2}$ (23)
@@
Figure 4: An instance of the randomized entangling ansatz (RAG) on $4$ qubits. It consists of an initial block of parameterized single-qubit rotations $3$ blocks of successive entangling operations. The entangling blocks themselves consist of a CZ gate and parameterized single-qubit rotation on the two involved qubits. The control and target for each two-qubit block are determined uniformly at random, i.e. the exact instance of an ansatz is defined by the random seed that was used for generating it. A more general version of this ansatz is detailed in Fig. 13.
@@
(b) Asymmetric depolarizing noise with noise strength $p=0.1$ and asymmetry $c=0.5$ following Eq. 41.
Figure 5: Distinguishability loss for different QEC codes on (a) depolarizing and (b) asymmetric depolarizing noise. All bars visualize the average-case loss $\mathcal{D}$ and worst-case loss $\overline{\mathcal{D}}$. Additionally, the two-design approximations $\mathcal{D}_{\mathcal{S}},\overline{\mathcal{D}}_{\mathcal{S}}$ are visualized as dotted lines if the approximation was not exact. For both noise models, the baseline of an unencoded qubit, the [[$5,1,3$]] perfect code [53], and VarQEC codes with $n=4$, $n=5$, and $n=7$ physical qubits are compared. All VarQEC ansätze are trained for $10$ epochs and contain $(n-1)\cdot(n-2)$ two-qubit blocks. Results on additional noise models can be found in Fig. 14, an extension to two data qubits is discussed in Fig. 16.
This theorem provides theoretical motivation for minimizing the distinguishability loss when designing encoding operations. As the bound assumes just a simple recovery operation, in practice, it is usually loose, and we can achieve an even better recovery fidelity. In the experimental section, we also empirically demonstrate that a lower distinguishability loss correlates with higher fidelity after recovery.
5 Empirical Evaluations
In this section, we present empirical evaluations of the proposed VarQEC method. We, for now, focus on training and validation in simulation, which we will complement with deployment of actual quantum devices in Section 6. The employed ansätze are detailed in Section D.1; a small exemplary instance is also shown in Fig. 4. Further implementation details of the developed VarQEC framework are discussed in Section D.2. We assess the performance of VarQEC codes under symmetric and asymmetric depolarizing noise in the main part, and extend our considerations to various other noise models in Appendix E. We analyze the stability and resilience of the learned encodings and compare the results with standard QEC codes.
@@ 5.1 Distinguishability Loss of Encodings
(a) Asymmetric depolarizing noise with a noise strength of $p=0.1$ and asymmetry $c$ as denoted on the $x$-axis. For $c<1$ the noise landscape is dominated by $Z$ errors, while for $c>1$ mainly $X$ and $Y$ errors occur, compare Eq. 41.
(b) Performance of noisy encoding operations on asymmetric depolarizing noise with a noise strength of $p=0.1$ and asymmetry $c=0.5$. In contrast to the other experiments, the gates in the encoding themselves are assumed to be non-ideal. To simulate this, correlated depolarizing noise of strength as denoted on the $x$-axis is added after each two-qubit operation, compare Eq. 47.
Figure 6: Stability and resilience of standard QEC and VarQEC codes in setups of (a) varying noise asymmetry and (b) non-ideal encoding operations. All data points represent the respective worst-case distinguishability loss $\overline{\mathcal{D}}$. The [[$5,1$]] VarQEC code is always the same static one trained on symmetric depolarizing noise with noise strength $p=0.1$ and under the assumption of ideal encoding operations.
@@
We begin by evaluating the distinguishability loss of different QEC and VarQEC codes when subjected to depolarizing noise. In addition, we simulate the same noise on an unencoded qubit as a baseline. We confirm these findings analytically in Section B.1, where the theoretical predictions perfectly align with the empirical results. Fig. 5(a) illustrates the distinguishability loss for this baseline, the [[$5,1,3$]] perfect code [53], and VarQEC codes with parameters [[$4,1$]], [[$5,1$]] and [[$7,1$]]. For all instances, the worst-case loss $\overline{\mathcal{D}}$, as well as the average-case loss $\mathcal{D}$, is displayed. Additionally, we indicate the two-design approximation values if they deviate from the ground-truth. We note that these values are themselves estimated with $1000$ Haar-random states, which is, however, sufficient for an accuracy of $10^{-4}$ as discussed in Section B.4. In this context, we have to remember that the [[$5,1,3$]] perfect code is the smallest code that can correct for arbitrary single-qubit errors [53]. Therefore, it is expected that it serves as a kind of lower bound for the specific setup of symmetric depolarizing noise. Indeed, the smaller VarQEC code with only $n=4$ physical qubits cannot improve the loss beyond the unencoded baseline of $\overline{\mathcal{D}}=0.133$. However, if we consider VarQEC models with $n\geq 5$ qubits, these perform on par with the loss of $\overline{\mathcal{D}}=1.106$ of the perfect QEC code. In that respect, we consider the slightly improved average-case performance of the [[$5,1$]] code with $\mathcal{D}=0.070$ only an artifact of the evaluation procedure, and not an actual performance improvement. We empirically confirm that the [[$5,1$]] VarQEC code exhibits a code distance of $d=3$, allowing for correcting single-qubit errors with a high fidelity, for details see Section B.3. Moreover, in Section 5.3 we demonstrate that these VarQEC codes allow for practically exact error correction on depolarizing noise. Overall, the learned encoding is on par with the optimal classical code for this setup.
@@
We extend our evaluation to asymmetric depolarizing noise in Fig. 5(b); the details are laid out in Section A.2. Summarized, for a total noise strength of $p=0.1$ and an asymmetry factor of $c=0.5$, the probability of an isolated phase-flip is $p_{z}\approx 0.085$, while the other two probabilities are $p_{x}=p_{y}\approx 0.0075$. In the presence of this asymmetry, the [[$5,1,3$]] perfect code can no longer be assumed to be optimal. Indeed, already the VarQEC code with $n=4$ qubits significantly improves the worst-case distinguishability loss. Moreover, the loss values decrease further when increasing the number of physical qubits to $n=5$ and $n=7$.
@@
The results demonstrate that the VarQEC procedure can produce codes that are competitive on noise structures where optimal QEC codes are known, and develop novel codes for more complex noise models. We strengthen this hypothesis by generating codes for other noise channels in Section E.1. Additionally, we extend our analysis to the setup of multiple encoded logical qubits, in particular $k=2$ in Section E.2
@@ 5.2 Stability and Fault Resilience
We further analyze the stability and resilience of the VarQEC codes and VarQEC procedure in general in Fig. 6. The experiments in Fig. 6(a) on varying asymmetry levels highlight the adaptability of the VarQEC procedure. For all setups from $c=0.5$, i.e. mostly phase-flips, to $c=2.0$, i.e. mostly bit-flips and combined bit- and phase-flips, all codes exhibit a distinguishability loss significantly beyond the unencoded baseline. The static version of the [[$5,1$]] VarQEC code (marked by [[$5,1$]]), i.e. the code learned on the symmetric setup $c=1.0$ from Fig. 5(a), outperforms the [[$5,1,3$]] perfect QEC code especially on setups with high asymmetries. Moreover, as one could expect, the improvement when training the VarQEC codes explicitly on the respective noise structure also increases with larger asymmetries. In this context, we refer to the codes noted by [[$5,1$]] as dynamically trained codes.
@@
In a more realistic scenario, one has to assume that the encoding operations themselves are non-ideal, which is typically considered within the context of fault-tolerant quantum computing [86]. However, this fault-tolerance analysis typically requires some highly formalized understanding of the involved encoding circuit structures, which has not been established yet for the VarQEC codes. Therefore, we approach this topic from a more empirical perspective, and therefore dub this analysis as the fault-resilience of the encodings. In Fig. 6(b), we do this by adding correlated depolarizing noise (see Eq. 47) of increasing strength after each two-qubit operation in the encoding. This is motivated by the observation that on current hardware systems, most error-prone operations are these involving multiple qubits. While single-qubit fidelity is already quantified around $99.99\%$, two-qubit gate fidelity is just recently approaching the infamous three 9’s [63, 29]. We do not claim that these gate fidelities can be directly related to the noise strength in Fig. 6(b). However, one can see that for a two-qubit gate fidelity beyond a certain threshold, the VarQEC encodings are reasonably fault-resilient. Moreover, for a wide range of error rates, they outperform the (also not fault-tolerant) encoding of the [[$5,1,3$]] perfect code, which only gradually catches up due to shallower circuit depth. Simulating the two-qubit errors already during the training procedure and therefore dynamically tailoring the codes further improves the results. The successful deployment of VarQEC encodings on quantum hardware we conduct in Section 6 underlines these fault-resilience properties.
@@
In Section E.4 we analyze fault-resilience under the assumption of restricted connectivity, which further strengthens the case for the VarQEC encodings. It is left for future work to enrich the approach with guarantees on actual fault-tolerance, e.g. by sufficiently extending the learning-based approach to simultaneously learn flag-gadgets [22, 104].
@@
(b) Asymmetric depolarizing noise with noise strength $p=0.1$ and asymmetry $c=0.5$ following Eq. 41.
Figure 7: Fidelity loss for different QEC codes and respective recovery operations on (a) depolarizing and (b) asymmetric depolarizing noise. All bars visualize the average-case loss $\mathcal{F}$ and worst-case loss $\overline{\mathcal{F}}$. For both noise models, the baseline of an unencoded qubit and the [[$5,1,3$]] perfect code are visualized. Additionally, we compare three instances of VarQEC codes, where [[$5,1$]] employs a random encoding, [[$5,1$]]∗∗ employs an encoding that has converged to a locally optimal distinguishability loss, and the encoding of [[$5,1$]] has converged to a global optimum. The recovery operations are successively trained following Eq. 22. For the rightmost model, the encoding and recovery operations have been trained end-to-end with the QVECTOR approach [46] as described in Appendix C. The recovery ansätze are trained using the fidelity loss for $50$ epochs and contain $200$ two-qubit blocks. As we consider a single-recovery scenario, no separate recovery register was used, i.e. the recovery operation acts solely on the $n$-qubit encoded and disturbed state [46]. The distinguishability loss for all encodings is summarized in Table 1. Results on additional noise models can be found in Fig. 15.
Base- [[$5,1,3$]] [[$5,1$]] [[$5,1$]]∗∗ [[$5,1$]] [[$5,1$]]
@@
Table 1: Worst-case distinguishability loss $\overline{\mathcal{D}}$ exhibited by the encodings for the models in Fig. 7. For the baseline and [[$5,1,3$]] code, this is evaluated on the unencoded baseline and static encoding operation, respectively. For the VarQEC codes, this corresponds to an untrained encoding for [[$5,1$]], to a local optimum for [[$5,1$]]∗∗, and to a global optimum for [[$5,1$]]. For the [[$5,1$]] QVECTOR code, the same ansatz with $12$ two-qubit blocks as for the VarQEC codes is employed, but training was only implicitly conducted via the fidelity loss. Additionally, we again report the worst-case fidelity loss $\overline{\mathcal{F}}$ (third line) and the respective upper (second line) and lower bounds (fourth line) following Theorems 2 and 1 for the VarQEC models.
5.3 Fidelity Loss with Recovery
Finally, we analyze the achievable fidelities after recovery for the noise channels and models we already considered in Section 5.1. This is quantified using the worst-case and average-case fidelity loss from Eqs. 21 and 19, which is evaluated over a two-design. In Fig. 7 we compare the unencoded baseline and $3$ code types: First, the [[$5,1,3$]] perfect QEC code with its static encoding and recovery operation. Second, three versions of a VarQEC code with $n=5$ physical qubits, one denoted as [[$5,1$]] with a randomly initialized untrained encoding, one referred to as [[$5,1$]]∗∗ where the distinguishability loss has converged to a local optimum, and the [[$5,1$]] VarQEC code from Fig. 5 where the encoding produces a presumably globally optimal distinguishability loss. The recovery operation is trained successively following Eq. 22. Third, a code with $n=5$ physical qubits where the encoding is not pre-trained, but encoding and recovery operations are simultaneously learned with the QCVECTOR approach [46], for details refer to Appendix C. We note that as we consider the single-recovery scenario, we do not employ additional recovery qubits for the variational models [46]. We list the respective distinguishability losses of the encodings and the bounds on the fidelity for the VarQEC models following Theorems 2 and 1 in Table 1.
@@ 5.3 Fidelity Loss with Recovery
The results in Fig. 7(a) on symmetric depolarizing noise demonstrate that a lower distinguishability loss of the encoding correlates with a lower fidelity loss, i.e. a higher fidelity after recovery. The results on the untrained VarQEC encoding and the one converged to a local optimum are below or just at the unencoded baseline, which further underlines the importance of the training procedure. We highlight that the optimal [[$5,1$]] VarQEC code achieves, up to numerical accuracy, an identical fidelity of $1-0.053=0.947$ as the [[$5,1,3$]] perfect code, which underlines the potential code distance of $d=3$ of this code (see Section B.3) and the interpretation as practically exact error correction. The [[$5,1$]] QVECTOR code is not successful in improving the fidelity beyond the unencoded baseline of $1-0.067=0.933$. We can observe qualitatively similar results on the asymmetric depolarizing noise channel in Fig. 7(b). Moreover, here the [[$5,1$]] VarQEC code guarantees a worst-case fidelity of $1-0.047=0.953$, which significantly improves upon the [[$5,1,3$]] QEC code with a fidelity of only $1-0.070=0.930$. Again, the QVECTOR approach is not able to improve the fidelity beyond the unencoded baseline.
@@
Regarding the QVECTOR results, we want to note that we employed the same ansätze as for the best VarQEC codes. Consequently, the underlying models in principle are guaranteed to be expressive enough to also produce the same competitive results. However, for the considered setups, the end-to-end training procedure based purely on the fidelity loss is unable to produce sophisticated codes. The results in Table 1 indicate that this is mostly because the resulting encoding operations do not sufficiently protect the state information before application of the noise channel. This is also substantiated by experiments on additional noise channels in Section E.1. This does not contradict the results from the original QVECTOR work [46], as we consider more complex and stronger noise channels and employ ansätze with significantly fewer parameters.
@@
(a) The [[$4,1$]] VarQEC code is trained in simulation for a noise duration of $10.0\mu\text{s}$. We deploy the model on the QPU for delays varying from $2.5\mu\text{s}$ to $10.0\mu\text{s}$. The bar plots visualize the distribution of the worst-case distinguishability loss over multiple logical patches.
(b) VarQEC codes with increasing number of physical qubits, i.e. [[$3,1$]], [[$4,1$]], and [[$5,1$]] are trained in simulation for a noise duration of $10.0\mu\text{s}$. To account for additional sources of errors, like noise introduced during the encoding operation itself and measurement delays, the delay is reduced to $5.0\mu\text{s}$ for hardware deployment. The violin plots visualize the distribution of the worst-case distinguishability loss over multiple logical patches.
Figure 8: Experiment on the ibm_marrakesh device [79], which features a 156-qubit Heron r2 chip. We induce thermal relaxation noise by applying a delay on all wires for a specific duration, assuming a median $T_{1}=180\mu\text{s}$ and $T_{2}=120\mu\text{s}$. The VarQEC codes are trained in classical simulation assuming these noise hyperparameters. Evaluation is conducted on hardware simultaneously on $8$ logical patches distributed uniformly over the QPU. We successively apply cluster re-sampling to enhance the granularity of data points [16, 67]. Additional details on the experimental procedure and the reconstruction of the worst-case distinguishability loss $\overline{\mathcal{D}}$ via mitigated state tomography can be found in the appendix. The predictions of the distinguishability loss from the simulation can be found in Fig. 21(a).

### v2

@@ 4.3 Generating Recovery Operations
While the main focus of our paper is the encoding operation, it is also important to construct fitting recovery operations $U_{\mathrm{rec}}$ for practical error correction. Therefore, we adopt a learning-based method inspired by the QVECTOR procedure [50] to obtain optimal parameters $\Phi$ for a variational ansatz $U_{\mathrm{rec}}(\Phi)$:
@@
[…]
@@
While our training of recovery operations in equation˜26 is only slightly modified from the QVECTOR procedure [50], the overall VarQEC framework is substantially more general and should not be viewed as a merely incremental extension. In QVECTOR, encoding and recovery are optimized simultaneously with respect to an average-case fidelity loss, tightly coupling the learned encoding to a particular recovery ansatz. In contrast, VarQEC first learns an encoding by minimizing the distinguishability loss $\overline{\mathcal{D}}$ via its average-case two-design proxy $\mathcal{D}_{\mathcal{S}}$), which both are based on the trace distance. This procedure comes with analytic guarantees relating $\overline{\mathcal{D}}$ to the worst-case fidelity after recovery, and an approximate code distance, as discussed in section˜4.4. Only in a second, modular step, we train the recovery operation using a QVECTOR-like fidelity objective, so that QVECTOR effectively appears as one possible instantiation of the recovery stage on top of our novel encoding-centric objective. This separation enables using the same learned encodings with different coherent (or, in future work, potentially also measurement-based) recovery strategies, which is not supported by the original end-to-end QVECTOR formulation. Empirically, we find that this two-stage, distinguishability-driven approach yields encodings that match or surpass established codes and consistently outperform QVECTOR-style end-to-end training, where fidelity-only optimization often fails to converge to useful solutions (see section˜5.3).
@@ 4.4 Properties of Learned Codes
However, we find that certain instances of VarQEC codes can achieve exact error correction. In particular, we verify this by evaluating the distinguishability loss of the code under all possible weight-$w$ errors. If all of these error patterns yield zero loss, then the code is said to have code distance $d=w+1$. The empirical method is described in more detail in section˜B.3. In particular, this analysis reveals that VarQEC codes with parameters $((n\geq 5,2))$ trained on depolarizing noise exhibit a code distance of $d=3$, matching the $[[5,1,3]]$ perfect code [57]. However, we again want to emphasize that the objective of VarQEC is not to generate codes with a certain provable code distance, but to develop low-overhead codes tailored for the specific noise channel.
As further motivation for training with the distinguishability loss, we derive a theoretical relationship to the worst-case fidelity after recovery. We state this connection informally here, with the formal proof deferred to section˜B.2.
@@
The worst-case distinguishability loss $\overline{\mathcal{D}}$ under a noise channel $\mathcal{N}$ guarantees the existence of a recovery operation that achieves a worst-case fidelity loss bounded by:
$\overline{\mathcal{F}}(\mathcal{N})\leq 1-\left(1-\overline{\mathcal{D}}(\mathcal{N})\right)^{2}$ (27)
@@
This theorem provides theoretical motivation for minimizing the distinguishability loss when designing encoding operations. We note that the bound of equation˜27 is not tight, i.e. in practice one can achieve an even better recovery fidelity. In the experimental section, we also empirically demonstrate that a lower distinguishability loss correlates with higher fidelity after recovery.
5 Empirical Evaluations
@@
(b) Asymmetric depolarizing noise with noise strength $p=0.1$ and asymmetry $c=0.5$ following equation˜45.
Figure 5: Distinguishability loss for different QEC codes on (a) depolarizing and (b) asymmetric depolarizing noise. All bars visualize the average-case loss $\mathcal{D}$ and worst-case loss $\overline{\mathcal{D}}$. Additionally, the two-design approximations $\mathcal{D}_{\mathcal{S}},\overline{\mathcal{D}}_{\mathcal{S}}$ are visualized as dotted lines if the approximation was not exact. For both noise models, the baseline of an unencoded qubit, the $[[5,1,3]]$ perfect code [57], and VarQEC codes with $n=4$, $n=5$, and $n=7$ physical qubits are compared. All VarQEC ansätze are trained for $10$ epochs and contain $(n-1)\cdot(n-2)$ two-qubit blocks. Results on additional noise models can be found in figure˜14, an extension to two data qubits is discussed in figure˜16.
In this section, we present empirical evaluations of the proposed VarQEC method. We, for now, focus on training and validation in simulation, which we will complement with deployment of actual quantum devices in section˜6. The employed ansätze and the experimental setup are detailed in Appendix˜D. We assess the performance of VarQEC codes under symmetric and asymmetric depolarizing noise in the following, and defer considerations of various other noise models and setups with multiple logical qubits to Appendix˜E.
@@ 5.1 Distinguishability Loss of Encodings
We begin by evaluating the distinguishability loss of different QEC and VarQEC codes when subjected to depolarizing noise. In addition, we simulate the same noise on an unencoded qubit as a baseline (see section˜B.1 for analytical justification). Figure˜5(a) illustrates the distinguishability loss for this baseline, the $[[5,1,3]]$ perfect code [57], and VarQEC codes with parameters $((4,2))$, $((5,2))$ and $((7,2))$. For all instances, the worst-case loss $\overline{\mathcal{D}}$, as well as the average-case loss $\mathcal{D}$, is displayed. Additionally, we indicate the two-design approximation values if they deviate from the ground-truth. We note that these values are themselves estimated with $1000$ Haar-random states, which is, however, sufficient for an accuracy of $10^{-4}$ (see section˜B.4). In this context, we have to remember that the $[[5,1,3]]$ perfect code is the smallest code that can correct for arbitrary single-qubit errors [57]. Therefore, it is expected that it serves as a kind of lower bound for the specific setup of symmetric depolarizing noise. Indeed, the smaller VarQEC code with only $n=4$ physical qubits cannot improve the loss beyond the unencoded baseline of $\overline{\mathcal{D}}=0.133$. However, if we consider VarQEC models with $n\geq 5$ qubits, these perform on par with the loss of $\overline{\mathcal{D}}=1.106$ of the perfect QEC code. In that respect, we consider the slightly improved average-case performance of the $((5,2))$ code with $\mathcal{D}=0.070$ only an artifact of the evaluation procedure, and not an actual performance improvement. We empirically confirm that the $((5,2))$ VarQEC code exhibits a code distance of $d=3$, allowing for correcting single-qubit errors with a high fidelity (see section˜B.3). Moreover, in section˜5.3 we demonstrate that these VarQEC codes allow for practically exact error correction on depolarizing noise. Overall, the learned encoding is on par with the optimal classical code for this setup.
@@
We extend our evaluation to asymmetric depolarizing noise in figure˜5(b). This noise channel represents general Pauli noise in the sense of equation˜4, instantiated by a bias parameter $c=\frac{\log p_{Z}}{\log p_{X}}=\frac{\log p_{Z}}{\log p_{Y}}$ [78]. For a total noise strength of $p=0.1$ and an asymmetry factor of $c=0.5$, this yields the probability of an isolated phase-flip is $p_{z}\approx 0.085$, while the other two probabilities are $p_{x}=p_{y}\approx 0.0075$. In the presence of this asymmetry, the $[[5,1,3]]$ perfect code can no longer be assumed to be optimal. Indeed, already the VarQEC code with $n=4$ qubits significantly improves the worst-case distinguishability loss. Moreover, the loss values decrease further when increasing the number of physical qubits to $n=5$ and $n=7$.
@@
The results demonstrate that the VarQEC procedure can produce codes that are competitive on noise structures where optimal QEC codes are known, and develop novel codes for more complex noise models.
@@ 5.2 Stability and Fault Resilience
(a) Asymmetric depolarizing noise with a noise strength of $p=0.1$ and asymmetry $c$ as denoted on the $x$-axis. For $c<1$ the noise landscape is dominated by $Z$ errors, while for $c>1$ mainly $X$ and $Y$ errors occur, compare equation˜45.
(b) Performance of noisy encoding operations on asymmetric depolarizing noise with a noise strength of $p=0.1$ and asymmetry $c=0.5$. In contrast to the other experiments, the gates in the encoding themselves are assumed to be non-ideal. To simulate this, correlated depolarizing noise of strength as denoted on the $x$-axis is added after each two-qubit operation, compare equation˜51.
Figure 6: Stability and resilience of standard QEC and VarQEC codes in setups of (a) varying noise asymmetry and (b) non-ideal encoding operations. All data points represent the respective worst-case distinguishability loss $\overline{\mathcal{D}}$. The $((5,2))^{*}$ VarQEC code is always the same static one trained on symmetric depolarizing noise with noise strength $p=0.1$ and under the assumption of ideal encoding operations.
@@
We further analyze the stability and resilience of the VarQEC codes and VarQEC procedure in general in figure˜6. The experiments in figure˜6(a) on varying asymmetry levels highlight the adaptability of the VarQEC procedure. For all setups from $c=0.5$, i.e. mostly phase-flips, to $c=2.0$, i.e. mostly bit-flips and combined bit- and phase-flips, all codes exhibit a distinguishability loss significantly beyond the unencoded baseline. The static version of the $((5,2))$ VarQEC code (marked by $((5,2)^{*}$), i.e. the code learned on the symmetric setup $c=1.0$ from figure˜5(a), outperforms the $[[5,1,3]]$ perfect QEC code especially on setups with high asymmetries. Moreover, as one could expect, the improvement when training the VarQEC codes explicitly on the respective noise structure also increases with larger asymmetries. In this context, we refer to the codes noted by $((5,2))$ as dynamically trained codes.
@@
In a more realistic scenario, one has to assume that the encoding operations themselves are non-ideal, which is typically considered within the context of fault-tolerant quantum computing [93]. However, this fault-tolerance analysis typically requires some highly formalized understanding of the involved encoding circuit structures, which has not been established yet for the VarQEC codes. Therefore, we approach this topic from a more empirical perspective, and therefore dub this analysis as the fault-resilience of the encodings. In figure˜6(b), we do this by adding correlated depolarizing noise (see equation˜51) of increasing strength after each two-qubit operation in the encoding. This is motivated by the observation that on current hardware systems, most error-prone operations are these involving multiple qubits. While single-qubit fidelity is already quantified around $99.99\%$, two-qubit gate fidelity is just recently approaching the infamous three 9’s [68, 31]. We do not claim that these gate fidelities can be directly related to the noise strength in figure˜6(b). However, one can see that for a two-qubit gate fidelity beyond a certain threshold, the VarQEC encodings are reasonably fault-resilient. Moreover, for a wide range of error rates, they outperform the (also not fault-tolerant) encoding of the $[[5,1,3]]$ perfect code, which only gradually catches up due to shallower circuit depth. Simulating the two-qubit errors already during the training procedure and therefore dynamically tailoring the codes further improves the results. The successful deployment of VarQEC encodings on quantum hardware we conduct in section˜6 underlines these fault-resilience properties.
@@
In section˜E.4 we analyze fault-resilience under the assumption of restricted connectivity, which further strengthens the case for the VarQEC encodings. It is left for future work to enrich the approach with guarantees on actual fault-tolerance, e.g. by sufficiently extending the learning-based approach to simultaneously learn flag-gadgets [23, 114].
5.3 Fidelity Loss with Recovery
Finally, we analyze the achievable fidelities after recovery for the noise channels and models we already considered in section˜5.1. This is quantified using the worst-case and average-case fidelity loss from equations˜25 and 23, which is evaluated over a two-design. In figure˜7 we compare the unencoded baseline and $3$ code types: First, the $[[5,1,3]]$ perfect QEC code with its static encoding and recovery operation. Second, three versions of a VarQEC code with $n=5$ physical qubits, one denoted as $((5,2))^{*}$ with a randomly initialized untrained encoding, one referred to as $((5,2))^{**}$ where the distinguishability loss has converged to a local optimum, and the $((5,2))$ VarQEC code from figure˜5 where the encoding produces a presumably globally optimal distinguishability loss. The recovery operation is trained successively following equation˜26. Third, a code with $n=5$ physical qubits where the encoding is not pre-trained, but encoding and recovery operations are simultaneously learned with the QVECTOR approach [50], for details refer to Appendix˜C. We note that as we consider the single-recovery scenario, we do not employ additional recovery qubits for the variational models [50]. We list the respective distinguishability losses of the encodings and the bounds on the fidelity for the VarQEC models following Theorems˜2 and 1 in table˜1.
@@ 5.3 Fidelity Loss with Recovery
(b) Asymmetric depolarizing noise with noise strength $p=0.1$ and asymmetry $c=0.5$ following equation˜45.
Figure 7: Fidelity loss for different QEC codes and respective recovery operations on (a) depolarizing and (b) asymmetric depolarizing noise. All bars visualize the average-case loss $\mathcal{F}$ and worst-case loss $\overline{\mathcal{F}}$. For both noise models, the baseline of an unencoded qubit and the $[[5,1,3]]$ perfect code are visualized. Additionally, we compare three instances of VarQEC codes, where $((5,2))^{*}$ employs a random encoding, $((5,2))^{**}$ employs an encoding that has converged to a locally optimal distinguishability loss, and the encoding of $((5,2))$ has converged to a global optimum. The recovery operations are successively trained following equation˜26. For the rightmost model, the encoding and recovery operations have been trained end-to-end with the QVECTOR approach [50] as described in Appendix˜C. The recovery ansätze are trained using the fidelity loss for $50$ epochs and contain $200$ two-qubit blocks. As we consider a single-recovery scenario, no separate recovery register was used, i.e. the recovery operation acts solely on the $n$-qubit encoded and disturbed state [50]. The distinguishability loss for all encodings is summarized in table˜1. Results on additional noise models can be found in figure˜15.
Base- [[$5,1,3$]] [[$5,1$]] [[5,15,1]]∗∗ [[$5,1$]] [[$5,1$]]
@@
Table 1: Worst-case distinguishability loss $\overline{\mathcal{D}}$ exhibited by the encodings for the models in figure˜7. For the baseline and $[[5,1,3]]$ code, this is evaluated on the unencoded baseline and static encoding operation, respectively. For the VarQEC codes, this corresponds to an untrained encoding for $((5,2))^{*}$, to a local optimum for $((5,2))^{**}$, and to a global optimum for $((5,2))$. For the $((5,2))$ QVECTOR code, the same ansatz with $12$ two-qubit blocks as for the VarQEC codes is employed, but training was only implicitly conducted via the fidelity loss. Additionally, we again report the worst-case fidelity loss $\overline{\mathcal{F}}$ (third line) and the respective upper (second line) and lower bounds (fourth line) following Theorems˜2 and 1 for the VarQEC models.
@@
The results in figure˜7(a) on symmetric depolarizing noise demonstrate that a lower distinguishability loss of the encoding correlates with a lower fidelity loss, i.e. a higher fidelity after recovery. The results on the untrained VarQEC encoding and the one converged to a local optimum are below or just at the unencoded baseline, which further underlines the importance of the training procedure. We highlight that the optimal $((5,2))$ VarQEC code achieves, up to numerical accuracy, an identical fidelity of 10.053=0.9471-0.053=0.947 as the $[[5,1,3]]$ perfect code, which underlines the potential code distance of $d=3$ of this code and the interpretation as practically exact error correction. The $((5,2))$ QVECTOR code is not successful in improving the fidelity beyond the unencoded baseline of 10.067=0.9331-0.067=0.933. We can observe qualitatively similar results on the asymmetric depolarizing noise channel in figure˜7(b). Moreover, here the $((5,2))$ VarQEC code guarantees a worst-case fidelity of 10.047=0.9531-0.047=0.953, which significantly improves upon the $[[5,1,3]]$ QEC code with a fidelity of only 10.070=0.9301-0.070=0.930. Again, the QVECTOR approach is not able to improve the fidelity beyond the unencoded baseline.
@@
Regarding the QVECTOR results, we want to note that we employed the same ansätze as for the best VarQEC codes. Consequently, the underlying models in principle are guaranteed to be expressive enough to also produce the same competitive results. However, for the considered setups, the end-to-end training procedure based purely on the fidelity loss is unable to produce sophisticated codes. The results in table˜1 indicate that this is mostly because the resulting encoding operations do not sufficiently protect the state information before application of the noise channel. This does not contradict the results from the original QVECTOR work [50], as we consider more complex and stronger noise channels and employ ansätze with significantly fewer parameters.